If p - q = 5 and pq = 3, find the value of p² + q².


-----
I've successfully solved a similar problem:
r² + s² = 15 and rs = 5, find the positive value of r + s.
r² + s² = 15
rs = 5
2 ⋅ rs = 5 ⋅ 2
2rs = 10
r² + 2rs + s² = 25
(r +s)² = 25
r + s = 5

Answer by noname9900(1)

square both sides
p ^2 + q ^2 - 2pq = 25
p ^2 + q ^2 = 25 + 6
p ^2 + q ^2 = 31